# Explaining Context Length Scaling and Bounds for Language Models

### v2

@@ Abstract
Long Context Language Models have drawn great attention in the past few years. There has been work discussing the impact of long context on Language Model performance: some find that long irrelevant context could harm performance, while some experimentally summarize loss reduction by relevant long context as Scaling Laws. This calls for a more thorough understanding on how long context impact Language Modeling. In this work, we (1) propose a clean and effective theoretical framework on explaining the impact of context length to Language Modeling, from an Intrinsic Space perspective; and (2) conduct experiments on natural language and synthetic data, validating our proposed theoretical assumptions and deductions. Our theoretical framework can provide practical insights such as establishing that training dataset size dictates an optimal context length and bounds context length scaling for certain case. We hope our work may inspire new long context Language Models, as well as future work studying Physics for Language Models. Code for our experiments is available at this anonymous url: https://anonymous.4open.science/r/xxxxxx.
Deep Learning, Context Length, Language Modeling
@@ 1 Introduction
Because of the rapid development of capacity of Language Models and the importance of a long context length in tasks like reasoning, retrieval, etc, past years people have been attempting to extend the context length of Language Models. There have been a variety of methods on supporting long context Language Models(Su et al., 2023; Katharopoulos et al., 2020; Gu & Dao, 2024; Peng et al., 2023; Sun et al., 2024). A wide variety of work is proposed to discuss the impact of context length: some shows long irrelevant context would worsen performance for LMs(Xu et al., 2024; Levy et al., 2024); some shows long context would improve performance in a way summarized as Scaling Laws(Xiong et al., 2024); while work in other domains like time series shows long relevant context would hurt performance (Shi et al., 2024). This calls for a more thorough understanding of how context length affects Language Models’ performance..
Previously, theories have been proposed to explain the Scaling Laws with respect to the data set and the size of the model(Bahri et al., 2024; Sharma & Kaplan, 2020). However, these theories do not study how context length impact Language Modeling, thus they cannot contribute directly to the problem.
In this work, we propose a theory framework to discuss the impact of context length from an Intrinsic Space perspective. Starting with simple assumptions w.r.t. Intrinsic Space and Intrinsic Dimension, we come up with a clean and effective theoretical explanation of relationship between Cross Entropy Loss, Intrinsic Dimension and Context Length. We further use real language and synthetic data to validate our assumptions and deductions. Our main contributions include,
@@
Our theoretical framework can naturally derive phenomena that we have observed. For example, our theory derives that, for a certain amount of training data, as the context length increases, the neural network first would first behave like the Bayes Model (thus the loss decreases); then beyond a certain optimal context length, the gap between the trained model and the Bayes Model would increase, hence the validation loss would increase: this is experimentally verified, as shown in Figure 1,
We hope our work may inspire future work when it comes to explaining context impact and/or designing new long context Language Models. For example, we observe that longer contexts have less impact (or there is a ceiling on the information about the next token as context length approaches infinity), discouraging models with excessively large context lengths. On the other hand, these findings have the potential to provide guarantees for RNNs with finite hidden states in long-sequence modeling.
Figure 1: Upper figure: Validation Loss vs. Context Length, measured on subsets of OpenWebText dataset. We see for each training dataset size, there exists an optimal context length that minimizes pretraining validation loss, which increases with the dataset size (More details can be found in Section 3). Lower figure: similar results obtained on our Synthetic Dataset in Section 4.5. This proves the deduction of our theory.
Our phenomenological theory in the main paper (mainly presented in Section 2) based on Intrinsic Dimension is simpler and more intuitive, while it can also be explained by a more fundamental theory based on Entropy, as presented in Appendix C.
@@ 2 Assumptions, Deductions and Observations for Language Modeling
2.1 Preliminary: Loss Decomposition and Intrinsic Space
It is common in ML studies to decompose the loss into Bayes Risk (the minimum loss possible, achieved by the theoretically optimal Bayesian Model), and Approximation Loss (the loss measuring the ability of a trained model actually to approximate the Bayesian Model). Specifically for Cross-Entropy loss, we have (please refer to Appendix A for more details):
$H(P,Q_{l})=$ $R_{Bayes}+L_{Approximate}$ (1)
$=$ $H(P,P_{l})+L_{Approximate}(P_{l},Q_{l})$
Where $P=p(x_{0}|x_{-\infty:0})$ is the distribution of Natural Language (or our experimented dataset), $P_{l}=p(x_{0}|x_{-l:0})$ is the Bayesian Model for context length $l$ and $Q=q(x_{0}|x_{-l:0})$ is the learned Language Model. $H(P,P_{l})$ is the Bayes Risk of optimal model (the assumed ‘limit’ when we have infinite data points and model parameters) and $L_{Approximate}(D,N,l,\ldots)$ is the Approximation Loss.
Bayesian Models are typically related to the Intrinsic Space of data manifold in such a way: the middle-layer or last-layer data representation of a well-trained Bayesian Neural Network is often used to approximate the Intrinsic Space(Cheng et al., 2023). In Section 2.2 we discuss more how this bridges Bayes Risk with Intrinsic Space, and how context length impacts Bayes Risk.
@@ 2.1 Preliminary: Loss Decomposition and Intrinsic Space
We use a synthetic data set to prove concepts in Section 4.
@@ 2.2.1 Bayes Risk and Entropy in Intrinsic Space: derived from First Principles
From first-principles we assume that,
@@
  • Assumption 3. Each intrinsic dimension would add $s$ bits of information to the next-token prediction task, so there are $dim(l)*s$ bits of information that can be represented in $\text{space}_{l}$ for the next-token prediction. Note this does not mean these are the only information in the Intrinsic Space, hence $s$ can be small, or even smaller than $1$.
  • Assumption 4. KL-divergence for the Bayes ¡+Model of context length $l$, $P_{l}$, with Bayes Model of infinite context length, $P=P_{\infty}$, equal to $s*(dim(\infty)-dim(l))$.
We further assume that in intrinsic space the states are with equal probability. With these assumptions, we can derive $H(P,P_{l})$ with $dim(l)$ (please refer to Appendix A for more details on the definition of Cross-Entropy loss used in this work, and Appendix B for a detailed derivation):
$R_{Bayes}(l)$ $=H(P,P_{l})$ (2)
$=H(P)+D_{KL}(P,P_{l})$
$\approx H(P)+s*(dim(\infty)-dim(l))$
$=-s*dim(l)+H(P)+s*dim(\infty).$
@@ 2.2.2 Bayes Risk and Intrinsic Dimension: Experiment Measurement
We further use PCA as a metric to measure the Intrinsic Dimension of Dataset with respect to context length. Figure 3 provides the relative degradation of the eigenvalue in the feature space of LLaMa-3.1-8B, for the last token. We see that larger input length would indeed provide feature with lower degradation in the intrinsic space. Notably, when $15005<idx<$ 1500 the curves is similar to Zip-f distribution ($\log eig=C_{0}-C*\log idx$), and for $4000500<idx<$ 4000 it resembles exponential degradation ($\log eig=C_{0}-C*idx$). Naturally, w ‘Intrinsic Dimension’. However, these formulas are very similar around $idx\approx 1000$, hence cannot provide accurate estimations to the transformation index.
Instead, following previous practice, here we use some threshold to decide the Ine would like to use the transformation index of these two states as Itrinsic Dimension: $\max_{idx}rela\_eig(idx)\geq\text{threshold}$ is used as the measured Intrinsic Dimensionz. We further evaluate the validity of this method in Point 2 in Section 4.1, on a synthetic Dataset with synthetic data; and explain the linear measurement from an Entropy perspective in Appendix C. Notably, the threshold here is a hyperparameter which is set to constants like $1/20$ in some previous work(Aghajanyan et al., 2020), but we observe here that many thresholds would validate the linear correspondence of Cross Entropy vs. Intrinsic Dimension, which further enhance the robustness of our result. In detail, as shown in Figure 3 and Figure 4, we use thresholds from $0.002$ to $0.25$ to make our conclusion more robust.
@@
Figure 3: Relative Eigen Value for LLaMa-3.1-8B on a subset of OpenWebText, presented in different x-axis scales, with different context length visible to Language Model. Gray lines represent different thresholds we take to measure the intrinsic dimension of the current model.
Figure 4: Cross Entropy Loss vs. Measured Intrinsic Dimension, for LLaMa-3.1-8B on a subset of OpenWebText. Each line represents a certain threshold used to measure ID in the intrinsic space of the used LLM. Different Measurements would give ID values that are linear w.r.t. each other, and they are all linear w.r.t. CE loss.
@@
We use the concept of Intrinsic Dimension and Measured Intrinsic Dimension to simplify the deduction. Moreover, by using the perspective of $S=\log\Omega$ where $\Omega$ is the number of states in Intrinsic Space, similar results can be derived, and why the measured IDs are linear w.r.t. each other and CE loss can be (partially) explained. Please refer to Appendix C for further discussions.
@@ 2.3 Approximation loss with Context Length: an intrinsic dimension perspective
In previous work people have experimentally summarize the Scaling Laws (Kaplan et al., 2020; Hoffmann et al., 2022) as: $L_{Approx}(D)=C_{0}+A/D^{\alpha}$ for different context length. Previous work has succeeded in explaining this from an intrinsic space perspective, represented by (Sharma & Kaplan, 2022) as: $\alpha\approx c/d_{intrinsic}$ where $c=2\text{ or }4$ based on model property, and $d_{intrinsic}$ is the dimension of the intrinsic space of the dataset.
@@
This shows that longer context length would make it harder for the model to learn to approximate the Bayesian Model.
@@ 3 Deduction: Optimal Context Length and Training Dataset Size
$L_{CE}$ $=C_{0}+\frac{C}{l^{\gamma}}+\frac{A(l)}{D^{\alpha(l)}},$ (6)
@@
We find $\lim_{D\rightarrow 0}f=-\infty$ and $\lim_{D\rightarrow\infty}f=\infty$. This shows that for fixed $l$, no matter what $\partial_{l}A$ is, there exists some $D$ s.t. $\partial_{l}L_{CE}=0$.
We see that Bayes Risk decreases with $l$, Approximation Loss increases with $l$ but decreases with $D$; the balance between these two losses results in an optimal $l$ that increases with the optimal $D$.
@@
As shown both theoretically and experimentally, when training until overfitting on the training dataset and considering the perplexity on validation dataset, there does exist an optimal context length, beyond which even relevant long context would increase validation loss of pretraining Language Models. Such optimal context length would increase with training dataset size.
@@ 4.1 List of Points to prove
  • Point 1. Cross Entropy Loss is approximately linear with Intrinsic Dimension. Shown in Section 4.3.
  • Point 2. By measuring Eigen-value degradation in Intrinsic Space of well-trained models, one could obtain a valid measurement of the Intrinsic Dimension. Shown in Section 4.4.
  • Point 3. There exists an optimal context length for each training dataset size used, and such optimal context length increases with the amount of training dataset. Shown in Section 4.5.
@@ 4.2 Construction of Synthetic Data: the ‘position weighted multitask sparse parity’ dataset
Figure 6: An example of the ‘two needles in a haystack’ task, similar to those in (Levy et al., 2024). Left: the input to the Language Model, with key information and question visualized in blue. Right: perplexity of the answer token $\langle 8\rangle$ of LLaMa-3.1-8B (horizontal axis) vs. number of masked leftmost tokens (vertical axis). Although seeing both pieces of information are necessary to answer the question, perplexity rises dramatically only when the first piece of information is masked.
@@
Inspired by this concept in Figure 6 and the ‘multitask sparse parity dataset previously studied in (Michaud et al., 2024; Barak et al., 2022), we propose the ‘position-weighted multitask sparse parity dataset. In detail, each input consists of $T$ ‘control bits and $L$ ‘context bits, each bit lies in $\{0,1\}$. Each subtask takes xor on two certain bits in the context bits, and the answer to some sample is the answer of the only activated subtask, as shown in Figure 7. We set tasks to have the same probability in the training and testing datasets. To emphasize the importance of context in different positions, for task $t$, we ensure $l(t)=max(context\_bit_{1}(t),context\_bit_{2}(t))$ follows $l(t)\approx 20/(1-t/50)^{1/1.2}$. This ensures,
@@
Figure 7: An example of our synthetic data. The answer for Subtask 1,2,3 is $0\oplus 0=0,0\oplus 1=1$ and $1\oplus 1=0$ respectively, but since the thrid bit is $1$ for control bits, only Subtask 3 is activated and the final answer is 00. However, for a model of context length $7$, it cannot see the $9th$ bit required by subtask 3, making it unable to predict the answer correctly.
@@
Thus we have, for some model who presents $s$ bits of information (about the output logit) in one Intrinsic Dimension:
@@
  • ID vs. CL: $ID\approx f(CL,C,C_{0},\gamma)$,$C_{0}=-51.1\pm 1.0$, $C=-1.7*10^{3}\pm 0.3*10^{3}$, $\gamma=1.18\pm 0.06$, $R^{2}=0.9997$.
  • CE vs. CL: $ID\approx f(CL,C,C_{0},\gamma)$,$C_{0}=-0.015\pm 0.013$, $C=23.8\pm 4.3$, $\gamma=1.18\pm 0.06$, $R^{2}=0.9997$.
  • CE vs. ID: $CE\approx g(ID,k,b)$, $k=0.693\pm 1*10^{-5}$, $b\approx 0.0139\pm 4*10^{-7}$, $R^{2}=1-7*10^{-9}$.
As shown, we construct synthetic data such that $ID(l)=ID_{0}-C^{\prime}/l^{\gamma}$, and our measurements show $CE=C-C^{\prime}/l^{\gamma}$. More importantly, for the synthetic data example, Cross Entropy loss is almost perfectly linear with the Intrinsic Dimension as we defined previously. This validates Point 1: Cross Entropy Loss is approximately linear with Intrinsic Dimension; and we have also provided a construction to match the measured relationship $CE(l)\approx C_{0}+C/l^{\gamma}$.
@@
Figure 8: Model trained on the proposed synthetic dataset; $\oplus$ represents feature concatenation. Only the first $l$ bits are used as input to context MLP when the context length is set to $l$. We conduct PCA on the Context Feature to analyze the intrinsic dimension of input context bits for various context lengths.
After training the model on data with different context length, we obtain the context feature representation and conduct PCA on it to study the Intrinsic Space of this model, presented in Figure 9.
@@
Figure 9: Illustration of relative eigen value vs. index, for models trained on different context length. Vertical lines: $x_{l}=ID(l)$ and $x_{l}^{\prime}=ID(l+1)$. For example, a context length $27$ has $16$ subtasks visible, corresponding to $16$ bits in Intrinsic Space. Assuming $1$ dimension in Intrinsic Space represents $1$ bit, the leftmost purple rectangle drawn means a range of $y_{threshold}$ that would provide an accurate estimation of $ID(27)=16$ for context length $27$. Obviously, there exists $y^{*}$ that would provide an estimation of $ID$ for all context lengths, as shown in the figure.
@@
The linear relationship between CE Loss and Intrinsic Dimension can also be explained from an Entropy in Intrinsic Space perspective, which is discussed in Appendix C.

### v3

@@ Abstract
Long Context Language Models have drawn great attention in the past few years. There has been work discussing the impact of long context on Language Model performance: some find that long irrelevant context could harm performance, while some experimentally summarize loss reduction by relevant long context as Scaling Laws. This calls for a more thorough understanding on how long context impacts Language Modeling. In this work, we (1) propose a clean and effective theoretical framework for explaining the impact of context length on Language Modeling, from an Intrinsic Space perspective; and (2) conduct experiments on natural language and synthetic data, validating our proposed theoretical assumptions and deductions. Our theoretical framework can provide practical insights such as establishing that training dataset size dictates an optimal context length and bounds context length scaling for certain cases. We hope our work may inspire new long context Language Models, as well as future work studying Physics for Language Models. [^1] Code for experiments is available at: https://github.com/JingzheShi/NLPCtlScalingAndBounds.
@@ 1 Introduction
Because of the rapid development of the capacity of Language Models and the importance of a long context length in tasks like reasoning, retrieval, etc., in recent years, people have been attempting to extend the context length of Language Models. There have been a variety of methods for supporting long context Language Models[1, 2, 3, 4, 5]. A wide variety of work is proposed to discuss the impact of context length: some shows long irrelevant context would worsen performance for LMs[6, 7]; some shows long context would improve performance in a way summarized as Scaling Laws[8]; while work in other domains like time series shows long relevant context would hurt performance [9]. This calls for a more thorough understanding of how context length affects Language Models’ performance.
Previously, theories have been proposed to explain the Scaling Laws with respect to the data set and the size of the model[10, 11]. However, most theories do not study how context length impacts Scaling Laws for Language Modeling, thus they cannot contribute directly to the problem. [^2] We discuss more about previous work in Appendix G.
In this work, we propose a theory framework to discuss the impact of context length from an Intrinsic Space perspective. Starting with simple assumptions w.r.t. Intrinsic Space and Intrinsic Dimension, we come up with a clean and effective theoretical explanation of the relationship between Cross Entropy Loss, Intrinsic Dimension and Context Length. We also use real language and synthetic data to validate our assumptions and deductions. Our main contributions include:
@@
Our theoretical framework (with formal definitions in Appendix C) can naturally derive phenomena that we have observed. For example, our theory derives that, for a certain amount of training data, as the context length increases, the neural network would first behave more similarly to the Bayes model (thus the loss decreases); while beyond a certain optimal context length, the gap between the trained model and Bayes Model would increase, hence validation loss would increase: this is experimentally verified in Figure 1,
We hope our work may inspire future work when it comes to explaining context impact and/or designing new long context Language Models. For example, we observe that longer contexts have less impact (or there is a ceiling on the information about the next token as context length approaches infinity), discouraging models with excessively large context lengths.
Figure 1: Left: Validation Loss Gap vs. Context Length, measured on subsets of OpenWebText dataset, where we subtract the minimum loss grouped by context length from each curve (please refer to Figure 13 for the original figure). We see for each training dataset size, there exists an optimal context length that minimizes pretraining validation loss, which increases with the dataset size (More details can be found in Section 3). Right: similar results obtained on our Synthetic Dataset in Section 4.5. This proves the deduction of our theory.
Our phenomenological theory in the main paper (mainly presented in Section 2) based on Intrinsic Dimension is simpler and more intuitive, while it can also be explained by a more fundamental theory based on Entropy, as presented in Appendix B. (We provide formal definitions of assumptions for both explanations in Appendix C.)
@@ 2 Assumptions, Deductions and Observations for Language Modeling
2.1 Preliminaries
2.1.1 Preliminary: Loss Decomposition
It is common in ML studies to decompose the loss into Bayes Risk (the minimum loss possible, achieved by the theoretically optimal Bayesian Model), and Approximation Loss (the loss measuring the ability of a trained model actually to approximate the Bayesian Model). Specifically for Cross-Entropy loss $H$, we have (please refer to Appendix A.1 for formal definitions and derivation details):
$H(P,Q_{l})=$ $R_{Bayes}+L_{Approx}$ (1)
$=$ $H(P,P_{l})+D_{KL}(P_{l},Q_{l})$
Where $P=p(x_{0}|x_{-\infty:0})$ is the distribution of Natural Language (or our experimented dataset), $P_{l}=p(x_{0}|x_{-l:0})$ is the Bayesian Model for context length $l$ and $Q_{l}=q(x_{0}|x_{-l:0})$ is the learned Language Model of context length $l$. $R_{Bayes}=H(P,P_{l})$ is the Bayes Risk of optimal model (the assumed ‘limit’ when we have infinite data points and model parameters) and $L_{Approx}=D_{KL}(P_{l},Q_{l})$ is the Approximation Loss, which can be affected by dataset size $D$, etc.
2.1.2 Preliminary: Intrinsic Space
In previous work[10, 12], as a common practice, the ‘Data Manifold’ is often defined as the middle feature representation of well-trained neural networks, and assumptions are made on this kind of mid-representation, with experiments to validate these assumptions. (Intrinsic Space is defined as the space where the Data Manifold lies.) We follow such practice in main paper for clarity.
Meanwhile, the Data Manifold can be more formally defined by a mapping from input data to some Intrinsic Space which satisfies a certain set of properties, and mid-representation of well-trained neural networks are assumed to have such properties, which can be experimentally validated. This is an equivalent yet more formal perspective. In Appendix C, we formally define the Intrinsic Space and derive related results in our work with such perspective for completeness.
2.1.3 Preliminary: Outlines
In Section 2.2 we discuss more how to bridge Bayes Risk with Intrinsic Space, and how context length impacts Bayes Risk. We further provide an Entropy-based explanation in Appendix B.
@@ 2.1.3 Preliminary: Outlines
We use a synthetic data set to prove concepts for Intrinsic dimension based explanations in Section 4, and entropy based explanations in Appendix B.
@@ 2.2.1 Bayes Risk and Entropy in Intrinsic Space: derived from First Principles
We assume these assumptions hold for Intrinsic Space (please see formal definitions of Intrinsic Space in Appendix C),
@@
  • Assumption 3. Uniform Information Gain ($s$-bits for next token prediction per Intrinsic Dimension): Each intrinsic dimension would add $s$ bits of information to the next-token prediction task, so there are $dim(l)*s$ bits of information that can be represented in $\text{space}_{l}$ for the next-token prediction. This means the KL-divergence for the Bayes Model of context length $l$, $P_{l}$, with Bayes Model of infinite context length, $P=P_{\infty}$, is: $D_{KL}(P,P_{l})=s*(dim(\infty)-dim(l))$. Note this does not mean these are the only information in the Intrinsic Space, hence $s$ can be small, or even smaller than $1$.
With these assumptions, we can derive $H(P,P_{l})$ with $dim(l)$ (please refer to Appendix A.1 for more details on the definition of Cross-Entropy loss used in this work, and assumptions and Theorem 3 in Appendix C for a detailed derivation):
$R_{Bayes}$ $=H(P,P_{l})$ (2)
$=-s*dim(l)+Const$
@@ 2.2.2 Bayes Risk and Intrinsic Dimension: Experiment Measurement
We further use PCA as a metric to measure the Intrinsic Dimension of Dataset with respect to context length. Figure 3 provides the relative degradation of the eigenvalue in the feature space of LLaMa-3.1-8B, for the last token. We see that larger input length would indeed provide feature with lower degradation in the intrinsic space. Notably, when $15005<idx<$ 1500 the curves is similar to Zip-f distribution ($\log eig=C_{0}-C*\log idx$), and for $4000500<idx<$ 4000 it resembles exponential degradation ($\log eig=C_{0}-C*idx$). [^3] Naturally, this would be a practical measure for the ‘Intrinsic Dimension’. However, these formulas are very similar around $idx\approx 1000$, hence cannot provide accurate estimations to the transformation index.
Instead, following previous practice, here we use some threshold to decide the Ine would like to use the transformation index of these two states as Itrinsic Dimension: $\max_{idx}rela\_eig(idx)\geq\text{threshold}$ is used as the measured Intrinsic Dimensionz. We further evaluate the validity of this method in Point 2 in Section 4.1, on a synthetic Dataset with synthetic data; and explain the linear measurement from an Entropy perspective in Appendix B. Notably, the threshold here is a hyperparameter which is set to constants in previous work(e.g.$1/20$ in [13]), but we observe here that many thresholds would validate the linear correspondence of Cross Entropy vs. Intrinsic Dimension, which further enhance the robustness of our result. In detail, as shown in Figure 3, we use thresholds from $0.002$ to $0.25$.
@@
Figure 3: Left figures: Relative Eigen Value for LLaMa-3.1-8B on a subset of OpenWebText, presented in different x-axis scales, with different context length visible to Language Model. Gray lines represent different thresholds we take to measure the intrinsic dimension of the current model. Right figures: Cross Entropy Loss vs. Measured Intrinsic Dimension. Each line represents a certain threshold used to measure ID in the intrinsic space of the used LLM. Different Measurements would give ID values that are linear w.r.t. each other, and they are all linear w.r.t. CE loss.
@@
We use the concept of Intrinsic Dimension and Measured Intrinsic Dimension to simplify the deduction. Moreover, by using the perspective of $S=\log\Omega$ where $\Omega$ is the number of states in Intrinsic Space, similar results can be derived, and why the measured IDs are linear w.r.t. each other and CE loss can be (partially) explained. Please refer to Appendix B for further discussions.
@@ 2.3 Approximation loss with Context Length: an intrinsic dimension perspective
Previous work experimentally summarizes the Scaling Laws [14, 15] as: $L_{Approx}(D)=C_{0}+A/D^{\alpha}$ for different dataset size $D$. Previous work has succeeded in explaining this from an intrinsic space perspective, represented by [10, 16] as: $\alpha\approx c/dim$, and $dim$ is the dimension of the data manifold of the data and the model, where a uniform distribution in Intrinsic Space is assumed. We derive this rigorously from weaker assumptions in Theorem 1,2 in Appendix C.2.
@@
This shows longer context length would make it harder for the model to learn to approximate the Bayes Model.
@@ 3 Deduction: Optimal Context Length and Training Dataset Size
$H(P,Q_{l})$ $=C_{0}+\frac{C}{l^{\gamma}}+\frac{A(l)}{D^{\alpha(l)}},$ (6)
@@
As shown, $\lim_{D\rightarrow 0}f=-\infty$ and $\lim_{D\rightarrow\infty}f=\infty$. This shows that for fixed $l$, no matter what $\partial_{l}A$ is, there exists some $D$ s.t. $\partial_{l}L_{CE}=0$.
Bayes Risk decreases with $l$, while Approximation Loss increases with $l$ but decreases with $D$; the balance between these two losses results in an optimal $l$ that increases with the optimal $D$.
@@
As shown both theoretically and experimentally, there does exist an optimal context length, beyond which even relevant long context would increase validation loss of pretraining Language Models. Such optimal context length would increase with training dataset size.
@@ 4.1 List of Points to prove
  • Point 1. Cross Entropy Loss is approximately linear with Intrinsic Dimension (Assumption 3 in Section 2.2.1). Shown in Section 4.3.
  • Point 2. By measuring Eigen-value degradation in Intrinsic Space of well-trained models, one could obtain a valid measurement of the Intrinsic Dimension (Section 2.2.2). Shown in Section 4.4.
  • Point 3. There exists an optimal context length for each training dataset size used, and such optimal context length increases with the amount of training dataset (Section 3). Shown in Section 4.5.
@@ 4.2 Construction of Synthetic Data: the ‘position weighted multitask sparse parity’ dataset
Figure 4: Left: An example of the ‘two needles in a haystack’ task, similar to those in [7]. The text part is the input to the Language Model, with key information and question visualized in blue; the figure part shows perplexity of the answer token $\langle 8\rangle$ of LLaMa-3.1-8B (horizontal) vs. number of masked leftmost tokens (vertical). Although seeing both pieces of information are necessary to answer the question, perplexity rises dramatically only when the first piece of information is masked. Right: An example of our synthetic data. The answer for Subtask 1,2,3 is $0\oplus 0=0,0\oplus 1=1$ and $1\oplus 1=0$ respectively, but since the thrid bit is $1$ for control bits, only Subtask 3 is activated and the final answer is 00. However, for a model of context length $7$, it cannot see the $9th$ bit required by subtask 3, making it unable to predict the answer correctly.
@@
Inspired by this concept in Figure 4 and the ‘multitask sparse parity dataset previously studied in [17, 18], we propose the ‘position-weighted multitask sparse parity dataset. In detail, each input consists of $T$ ‘control bits and $L$ ‘context bits, each bit lies in $\{0,1\}$. Each subtask takes xor on two certain bits in the context bits, and the answer to some sample is the answer of the only activated subtask, as shown in Figure 4. We set tasks to have the same probability in the training and testing datasets; we discuss cases where tasks have non-equal probability in Appendix B.6. To emphasize the importance of context in different positions, for task $t$, we ensure $l(t)=max(context\_bit_{1}(t),context\_bit_{2}(t))$ follows $l(t)\approx 20/(1-t/50)^{1/1.2}$. This ensures,
@@
Thus, for some model presenting $s$ bits of information (about output logit) in one Intrinsic Dimension:
@@
  • ID & CL: $ID\approx f(CL,C,C_{0},\gamma)$,$C_{0}=51.1\pm 1.0$, $C=1.7*10^{3}\pm 0.3*10^{3}$, $\gamma=1.18\pm 0.06$, $R^{2}=0.9997$.
  • CE & CL: $CE\approx f(CL,C,C_{0},\gamma)$,$C_{0}=-0.015\pm 0.013$, $C=-23.8\pm 4.3$, $\gamma=1.18\pm 0.06$, $R^{2}=0.9997$.
  • CE & ID: $CE\approx g(ID,k,b)$, $k=-0.693\pm 1*10^{-5}$, $b\approx 0.0139\pm 4*10^{-7}$, $R^{2}=1-7*10^{-9}$.
As shown, we construct synthetic data such that $ID(l)=ID_{0}-C^{\prime}/l^{\gamma}$, and our measurements show $CE=C+C^{\prime}/l^{\gamma}$. More importantly, for the synthetic data example, Cross Entropy loss is almost perfectly linear with the Intrinsic Dimension as we defined previously. This validates Point 1: Cross Entropy Loss is approximately linear with Intrinsic Dimension; and we have also provided a construction to match the measured relationship $CE(l)\approx C_{0}+C/l^{\gamma}$.
@@
Figure 5: Model trained on the proposed synthetic dataset; $\oplus$ represents feature concatenation. Only the first $l$ bits are used as input to context MLP when the context length is set to $l$. We conduct PCA on Context Feature to analyze the intrinsic dimension of input context bits for various context lengths.
After training the model on data with different context length, we conduct PCA on the obtain context feature representation to study the Intrinsic Space of this model, presented in Figure 6.
@@
Figure 6: Relative eigen value vs. index, for models trained on different context length. Vertical lines: $x_{l}=ID(l)$ and $x_{l}^{\prime}=ID(l+1)$. For example, a context length $27$ has $16$ subtasks visible, corresponding to $16$ bits in Intrinsic Space. Assuming $1$ dimension in Intrinsic Space represents $1$ bit, the leftmost purple rectangle drawn means a range of $y_{threshold}$ that would provide an accurate estimation of $ID(27)=16$ for context length $27$. There exists $y^{*}$ that would provide an estimation of $ID$ for all context lengths, as shown in the figure.
@@
The linear relationship between CE Loss and Intrinsic Dimension can also be explained from an Entropy in Intrinsic Space perspective, which is discussed in Appendix B.